OpenStudy (anonymous):

how weight changes depending on the acceleration due to gravity on jupiter earth and mars

OpenStudy (anonymous):

weight is directly proportional to acceleration due to gravit

OpenStudy (anonymous):

W = m g, but g will vary from planet to planet and even from one distance to another.

OpenStudy (jfraser):

the acceleration due to gravity depends directly on the mass of the planet and its radius, not the mass of the object on the planet. It's derived from Newton's second law and the Universal Gravitation:\[F_G = G*\frac{M_1*m_2}{r^2} = m_2*a_G\] Newton's 2nd Law and Universal Gravitation Law put together\[\cancel{m_2}*a_G = G*\frac{M_1*\cancel{m_2}}{r^2}\]so \[a_G = \frac{G*M_1}{r^2}\]where G is the universal gravitation constant, \(M_1\) is the mass of the planet, and \(r\) is the radius of the planet (assuming you're standing on the surface of the planet)